Topic: Rate Equations

Question 1

Twi compounds, A and B , react in the following way:

A + 3B —> AB3.

using the data below, derive the rate equation for this reaction and explain your reasoning.(4 marks)

b) Use the experimental data provided below to determine the rate constant. ( 3 marks)

data:

  • Experiment 1 ,Initial (A) ( mol dm-3) 0.100 , Initial (B) (mol dm3 ) 0.100 , Initial rate of formation of AB (mol dm-3)0.002.
  • Experiment 2, Initial (A) (mol dm-3) 0.100 , Initial (B) ( mol dm-3) 0.200, Initial rate of formation of AB3 ( mol dm-3) 0.008.
  • Experiment 3 , Initial (A) (mol dm-3) 0.100 , Initial (B) (mol dm-3) 0.300, Initial rate of formation of AB3 (mol dm-3) 0.018.
  • Experiment 4, Initial (A) ( mol dm-3) 0.200 , Initial (B) ( mol dm-3) 0.100 , Initial rate of formation of AB ( mol dm-3) 0.004.
  • Experiment 5, Initial (A) (mol dm-3) 0.300 , Initial (B) (mol dm-3) 0.100 , Initial rate of formation of AB (mol dm-3) 0.006.
1
Expert's answer
2022-01-18T03:01:03-0500

Let the rate be given by r = k[A]x[B]y


1) 0.002 = k(0.100)x(0.100)y

2) 0.008 = k(0.100)x(0.200)y

3) 0.018 = k(0.100)x(0.300)y

4) 0.004 = k(0.200)x(0.100)y

5) 0.006 = k(0.300)x(0.100)y


From 1, 2 and 3


"\dfrac{0.018}{0.008 } = \dfrac{k(0.100)^x(0.300)^y}{k(0.100)^x(0.200)^y}"

"2.25 = (1.5)^y => y = 2"

OR

"\dfrac{0.008}{0.002} = \dfrac{k(0.100)^x(0.200)^y}{k(0.100)^x(0.100)^y}"

"4 = (2)^y => y = 2"


From 1, 4 and 5


"\dfrac{0.004}{0.002} = \dfrac{k(0.200)^x(0.100)^y}{k(0.100)^x(0.100)^y}"

"2 = (2)^x => x = 1"

OR

"\dfrac{0.006}{0.002} = \dfrac{k(0.300)^x(0.100)^y}{k(0.100)^x(0.100)^y}"

"3 = (3)^x => x = 1"


The rate equation for this reaction r = k[A][B]2

From experiment 1 the rate constant

"k=\dfrac{0.002\ mol\ dm^{-3}\ s^{-1}}{0.100\ mol\ dm^{-3}\ *(0.100\ mol\ dm^{-3})^2}=2\ dm^6\ mol^{-2}\ s^{-1}"